# Kengo Nakamura, Masaaki Nishino, and Shuhei Denzumi NTT Communication Science Laboratories, Kyoto, Japan

### v1

@@ Abstract
Zero-suppressed binary decision diagram (ZDD) is a data structure to represent a family of (sub)sets compactly, and it can be used as a succinct index for a family of sets. To build ZDD representing a desired family of sets, there are many transformation operations that take ZDDs as inputs and output ZDD representing the resultant family after performing operations such as set union and intersection. However, except for some basic operations, the worst-time complexity of taking such transformation on ZDDs has not been extensively studied, and some contradictory statements about it have arisen in the literature. In this paper, we show that many transformation operations on ZDDs cannot be performed in worst-case polynomial time with respect to the size of input ZDDs. This refutes some of the folklore circulated in past literature and resolves an open problem raised by Knuth. Our results are stronger in that such blow-up of computational time occurs even when the ordering, which has a significant impact on the efficiency of treating ZDDs, is reasonable.
@@ 1 Introduction
To alleviate this issue, we can use zero-suppressed binary decision diagram (ZDD[7], which is a variant of the well-known binary decision diagram (BDD) [2]. ZDD can represent a family of sets compactly. Moreover, it supports many queries about the represented family of sets, e.g., counting the number of sets and performing linear optimization over the family. Thus, ZDD can be used as a succinct index for a family of sets once the corresponding ZDD is built.
@@
ZDDs also support a number of transformation operations. For example, when we have two ZDDs representing two families of sets, we can construct a ZDD representing the set union of them without extracting each set from the input families. Using such operations, we can construct a ZDD representing the desired family of sets. By collecting such transformation operations, Minato [8] considered an algebraic system called unate cube set algebra, whose element is a family of sets. After that, many operations were introduced, and now the system is widely called family algebra, whose name was given by Knuth [6]. With the algorithms performing operations on ZDDs, every operation in the family algebra provides a useful way to construct a ZDD representing the desired family of sets in many applications.
@@
However, the time complexity of performing operations on ZDDs has not been well studied, except for basic set operations. This is because some operations require complicated recursion procedures that make complexity analysis difficult. In particular, revealing worst-case time complexity is important to us. If the worst-case time complexity is large, it takes an unexpectedly long time to carry out a single operation on ZDDs for certain kinds of input. If so, we should pay attention to the possibility of such input when we use ZDDs as a way to implement the manipulation of families of sets. Therefore, we investigated the worst-case time complexity of executing a single family algebra operation on ZDDs.
1.1 Related Works
@@
Since the invention of ZDD [7], many family algebra operations have been proposed. Table 1 lists the basic operations of family algebra. As related works, we first describe the origins of these operations.
@@
The first three operations, union, intersection, and difference, are the most fundamental set operations set described by Minato [7]. The join, quotient, and remainder operations appeared in Minato’s next paper [8], where the join operation is called “product” because a join can be considered to be the multiplication of two families when we view the union operation as an addition operation. These operations are peculiar to the families of sets and also fundamental in defining other family algebra operations. Later, the disjoint join and joint join operations were proposed by Kawahara et al. [5] through an extension of the join; their usage is intended to implicitly enumerate all of the subgraphs having a particular shape.
@@
Restrict and permit operations were originally proposed by Coudert et al. [4], where they are called SupSet and SubSet and used for solving set cover problems or performing logic circuit minimization. The names “restrict” and “permit” come from a study by Okuno et al. [10]. Later, nonsuperset, nonsubset, maximal, and minimal operations were introduced by Coudert [3] to solve various optimization problems on graphs. Furthermore, meet, delta, and minimal hitting set operations were introduced by Knuth [6] to solve various graph problems. In this way, the family algebra operations have been gradually developed.
Table 1: List of Operations on Family Algebra
Operation Definition
Union $\mathcal{F}\cup\mathcal{G}$$\{S\mid S\in\mathcal{F}\vee S\in\mathcal{G}\}$
Intersection $\mathcal{F}\cap\mathcal{G}$$\{S\mid S\in\mathcal{F}\wedge S\in\mathcal{G}\}$
Difference $\mathcal{F}\setminus\mathcal{G}$$\{S\mid S\in\mathcal{F}\wedge S\notin\mathcal{G}\}$
Join $\mathcal{F}\sqcup\mathcal{G}$$\{F\cup G\mid F\in\mathcal{F},G\in\mathcal{G}\}$
Disjoint join ˙$\mathcal{F}\operatorname{\dot{\bowtie}}\mathcal{G}$$\{F\cup G\mid F\in\mathcal{F},G\in\mathcal{G},F\cap G=\emptyset\}$
Joint join ^$\mathcal{F}\operatorname{\hat{\bowtie}}\mathcal{G}$$\{F\cup G\mid F\in\mathcal{F},G\in\mathcal{G},F\cap G\neq\emptyset\}$
Meet $\mathcal{F}\sqcap\mathcal{G}$$\{F\cap G\mid F\in\mathcal{F},G\in\mathcal{G}\}$
Delta $\mathcal{F}\operatorname{\boxplus}\mathcal{G}$$\{F\oplus G\mid F\in\mathcal{F},G\in\mathcal{G}\}$
Quotient $\mathcal{F}\operatorname{/}\mathcal{G}$$\{S\mid\forall G\in\mathcal{G},S\cup G\in\mathcal{F}\wedge S\cap G=\emptyset\}$
Remainder %$\mathcal{F}\operatorname{\%}\mathcal{G}$ $\mathcal{F}\setminus(\mathcal{G}\sqcup(\mathcal{F}\operatorname{/}\mathcal{G}))$
Restrict $\mathcal{F}\bigtriangleup\mathcal{G}$$\{F\in\mathcal{F}\mid\exists G\in\mathcal{G}\ \text{s.t.}\ G\subseteq F\}$
Permit $\mathcal{F}\oslash\mathcal{G}$$\{F\in\mathcal{F}\mid\exists G\in\mathcal{G}\ \text{s.t.}\ F\subseteq G\}$
Nonsuperset $\mathcal{F}\operatorname{\searrow}\mathcal{G}$$\{F\in\mathcal{F}\mid\forall G\in\mathcal{G},G\nsubseteq F\}$
Nonsubset $\mathcal{F}\operatorname{\nearrow}\mathcal{G}$$\{F\in\mathcal{F}\mid\forall G\in\mathcal{G},F\nsubseteq G\}$
@@
Maximal $\mathcal{F}^{\uparrow}$$\{F\in\mathcal{F}\mid\forall F^{\prime}\in\mathcal{F},F\subseteq F^{\prime}\Rightarrow F=F^{\prime}\}$
@@
Minimal $\mathcal{F}^{\downarrow}$$\{F\in\mathcal{F}\mid\forall F^{\prime}\in\mathcal{F},F^{\prime}\subseteq F\Rightarrow F=F^{\prime}\}$
@@
Minimal hitting set $\mathcal{F}^{\sharp}$$\{S\mid\forall F\in\mathcal{F},S\cap F\neq\emptyset\}^{\downarrow}$
@@
Compared to the operations themselves, the time complexity of performing them on ZDDs has not been well investigated. Minato [7] proved that the first three operations in Table 1 can be performed in polynomial time with respect to the size of input ZDDs. However, the complexity of a join operation, the most basic one among the rest, has not been fully clarified. Knuth [6] claimed in the answer section of exercises that join, as well as meet and delta, can be performed in worst-case polynomial time, but this claim lacks proof. Conversely, Kawahara et al. [5] suggested that join, as well as disjoint join and joint join, take worst-case exponential time, again without proof. In addition to those reports, Okuno et al. [10] claimed that restrict can be performed in polynomial time, but they used the unproven proposition that join can be performed in polynomial time. Furthermore, Knuth [6] stated that the worst-case complexity of the quotient operation was an open problem.
Note that the average complexity of carrying out operations on ZDDs was empirically confirmed in some cases. For example, Minato [9] stated that the complexity of join is empirically polynomial with respect to the input ZDD size in many cases. Knuth [6] stated that join, meet, delta, quotient, and remainder operations are reasonably fast in many empirical cases. However, even the average complexity has not been rigorously proved.
@@ 1.2 Our Contribution
In this paper, we prove that for the operations in Table 1 aside from union, intersection, and difference, there exist polynomial-sized ZDDs such that after taking the operation, the ZDD size becomes exponential. For example, for the join operation, we prove that there exist sequences of families of sets $\{\mathcal{F}_{m}\}$ and $\{\mathcal{G}_{m}\}$ such that the ZDD sizes representing $\mathcal{F}_{m}$ and $\mathcal{G}_{m}$ are polynomial in $m$, while the ZDD size representing $\mathcal{F}_{m}\sqcup\mathcal{G}_{m}$ is exponential in $m$. This result implies that these operations cannot be performed in worst-case polynomial time with respect to the size of input ZDDs. Thus, we refute the statement raised by Knuth [6] and Okuno et al. [10] that join, meet, delta, and restrict can be performed in worst-case polynomial time. We also resolve the open problem, i.e., the worst-case complexity of the quotient operation, raised by Knuth [6].
@@
Our result is stronger in that the resultant ZDD’s size remains exponential for any order of elements. ZDD structures follow a total order of the elements in the base set, and it is known that this element order has a significant impact on the ZDD size. However, we also prove that for the sequences used in proving the above, the resultant ZDD’s size is exponential in $m$ regardless of the order of elements. This suggests that we cannot shrink the ZDD size after taking an operation by managing the element order. Some famous BDD manipulation packages such as CUDD [11] implemented dynamic reordering, the reordering of elements after executing operations to shrink the ZDD size and thus increase the efficiency of BDD/ZDD manipulations. Nevertheless, our results suggest that the worst-case complexity of carrying out operations cannot be polynomial, even if we employ reordering.
@@
Note that this follows the research line of Bollig [1] as follows. Yoshinaka et al. [12] refuted Bryant’s conjecture, which is about the complexity of performing operations on BDDs, but their counterexample was somewhat weak in that the order of elements they used was unfavorable for BDD representations. Bollig [1] later resolved this issue by proposing simpler counterexamples. Similar to this, our results imply that the exponential blow-up in taking an operation on ZDDs occurs not only when the order of elements is unfavorable but also when it is reasonable.
@@ 2 Preliminaries
In this section, we formally define the structure and semantics of ZDDs. After that, we give a toy example for some family algebra operations. Finally, we briefly explain how to carry out family algebra operations on ZDDs.
@@ 2.1 Zero-suppressed Binary Decision Diagram
A zero-suppressed binary decision diagram (ZDD[7] is a rooted directed acyclic graph (DAG)-shaped data structure for representing a family of sets. First, we describe the structure of ZDD. ZDD $\mathtt{Z}$ consists of node set $\mathtt{N}$ and arc set $\mathtt{A}$, where the node set contains terminal nodes $\top$,$\bot$ , ⊥ and other internal nodes. Terminal nodes have no outgoing arcs, while every internal node has two outgoing edges called lo-arc and hi-arc. The nodes pointed by the lo-arc and the hi-arc outgoing from a node $\mathtt{n}$ are called lo-child and hi-child of $\mathtt{n}$, and they are denoted by $\mathsf{lo}(\mathtt{n})$ and $\mathsf{hi}(\mathtt{n})$, respectively. Every internal node $\mathtt{n}$ is associated with an element called label that is denoted by $\mathsf{lb}(\mathtt{n})$. ZDDs must follow the ordered property: Given a total order of elements $<$, the label of the parent node must precede that of the child node, i.e., $\mathsf{lb}(\mathtt{n})<\mathsf{lb}(\mathsf{lo}(\mathtt{n}))$ and $\mathsf{lb}(\mathtt{n})<\mathsf{lb}(\mathsf{hi}(\mathtt{n}))$ must hold for every internal node $\mathtt{n}$. Note that the child node is always allowed to be a terminal node. Finally, the size of a ZDD is defined by its number of nodes.
Next, we describe the semantics of ZDD. Given ZDD $\mathtt{Z}$, the family of sets represented by $\mathtt{Z}$ is defined recursively as follows.
Definition 1.
@@
2.2 Toy Examples for Family Algebra Operations
Before explaining how to perform family algebra operations on ZDDs, we first describe some examples of family algebra operations for the sake of improving readability. Let =$\mathcal{F}=\{\{x_{1},x_{3}\},\{x_{2},x_{4}\}\}$ and $\mathcal{G}=\{\{x_{1},x_{3}\},\{x_{2}\},\{x_{3},x_{4}\}\}$. Then, the union of them is $\mathcal{F}\cup\mathcal{G}=\{\{x_{1},x_{3}\},\{x_{2}\},\{x_{2},x_{4}\},\{x_{3},x_{4}\}\}$; it is simply the set union when viewing every set as an element. The join of them is $\mathcal{F}\sqcup\mathcal{G}=\{\{x_{1},x_{3}\},\{x_{1},x_{2},x_{3}\},\{x_{1},x_{3},x_{4}\},\{x_{1},x_{2},x_{3},x_{4}\},\{x_{2},x_{4}\},\{x_{2},x_{3},x_{4}\}\}$. For example, $\{x_{1},x_{3},x_{4}\}\in\mathcal{F}\sqcup\mathcal{G}$, since it is the union of $\{x_{1},x_{3}\}\in\mathcal{F}$ and $\{x_{3},x_{4}\}\in\mathcal{G}$. For the restrict and permit operations, we have $\mathcal{F}\bigtriangleup\mathcal{G}=\{\{x_{1},x_{3}\},\{x_{2},x_{4}\}\}$ and $\mathcal{F}\oslash\mathcal{G}=\{\{x_{1},x_{3}\}\}$. Here, $\{x_{2},x_{4}\}\in\mathcal{F}\bigtriangleup\mathcal{G}$, since $\{x_{2},x_{4}\}\supseteq\{x_{2}\}\in\mathcal{G}$, and $\{x_{2},x_{4}\}\notin\mathcal{F}\oslash\mathcal{G}$, since it is not a subset of any set in $\mathcal{G}$.

### v2

@@ Abstract
Binary decision diagram (BDD) and zero-suppressed binary decision diagram (ZDD) are data structures to represent a family of (sub)sets compactly, and it can be used as succinct indexes for a family of sets. To build BDD/ZDD representing a desired family of sets, there are many transformation operations that take BDDs/ZDDs as inputs and output BDD/ZDD representing the resultant family after performing operations such as set union and intersection. However, except for some basic operations, the worst-time complexity of taking such transformation on BDDs/ZDDs has not been extensively studied, and some contradictory statements about it have arisen in the literature. In this paper, we show that many transformation operations on BDDs/ZDDs, including all operations for families of sets that appear in Knuth’s book, cannot be performed in worst-case polynomial time in the size of input BDDs/ZDDs. This refutes some of the folklore circulated in past literature and resolves an open problem raised by Knuth. Our results are stronger in that such blow-up of computational time occurs even when the ordering, which has a significant impact on the efficiency of treating BDDs/ZDDs, is chosen arbitrarily.
keywords:
Binary decision diagrams, family of sets, family algebra
category:
[…]
@@ 1 Introduction
To alleviate this issue, we can use binary decision diagram (BDD[2] or zero-suppressed binary decision diagram (ZDD[14] that is a variant of BDD. BDD and ZDD are data structures that compactly represent a Boolean function and a family of sets, respectively. Since a Boolean function $f$ can be regarded as a family of sets by considering the set of assignments of input Boolean variables that evaluates $f$ to true, BDD can also be regarded as a succinct representation of a family of sets. Moreover, they support many queries about the represented family of sets, e.g., counting the number of sets and performing linear optimization over the family. Thus, BDD and ZDD can be used as succinct indexes for a family of sets.
@@
BDDs and ZDDs also support a number of transformation operations. For example, when we have two BDDs representing two families of sets, we can construct a BDD representing the set union of them without extracting each set from the input families. Using such operations, we can construct a BDD or a ZDD representing the desired family of sets. By collecting such transformation operations, Minato [15] considered an algebraic system called unate cube set algebra, whose element is a family of sets. After that, many operations were introduced, and now the system is widely called family algebra, whose name was given by Knuth [13]. With the algorithms performing operations on BDDs and ZDDs, every operation in the family algebra provides a useful way to construct a BDD or a ZDD representing the desired family of sets in many applications. Many of these operations have been implemented in standard BDD and ZDD manipulation packages [8, 18], and they are used in a wide range of applications, including formal verification of circuits [7, 10], analyses of power distribution networks [9, 19], and data mining [16].
@@
However, the complexity of performing family algebra operations on BDDs and ZDDs has not been well studied, except for basic set operations. This is because some operations require complicated recursion procedures that make complexity analysis difficult. In particular, revealing worst-case time complexity is important to us. If the worst-case time complexity is large, it takes an unexpectedly long time to carry out even a single operation for certain kinds of input. If so, we should pay attention to the possibility of such input when we use BDDs and ZDDs as a way to implement the manipulation of families of sets. Therefore, we investigated the worst-case time complexity of executing a single family algebra operation on BDDs and ZDDs. Since it is known that, as described later, the sizes of a BDD and a ZDD representing the same family of sets differ in only a linear factor, this paper mainly focused on the complexity of ZDDs. After that, we mention the complexity on BDDs.
1.1 Related Work
@@
Since the invention of ZDD [14], many family algebra operations have been proposed. Table 1 lists basic operations. As related work, we first describe the origins of these operations.
@@
The first four operations in Table 1 are the most fundamental set operations set described by Minato [14]. The join, quotient, and remainder operations appeared in Minato’s next paper [15], where the join operation is called “product” because a join can be considered to be the multiplication of two families when we view the union operation as an addition operation. These operations are peculiar to the families of sets and also fundamental in defining other family algebra operations. Later, the disjoint join and joint join operations were proposed by Kawahara et al. [12] through an extension of the join; their usage is to implicitly enumerate all of the subgraphs having a particular shape.
@@
Restrict and permit operations were originally proposed by Coudert et al. [5], where they were called SupSet and SubSet and used for solving set cover problems or performing logic circuit minimization. The names “restrict” and “permit” come from a study by Okuno et al. [17]. Later, nonsuperset, nonsubset, maximal, and minimal operations were introduced by Coudert [4] to solve various optimization problems on graphs. Furthermore, meet, delta, minimal hitting set, and closure operations were introduced by Knuth [13, §7.1.4 Ex.203,236,243] to solve various graph problems. Table 1 contains all of the transformation operations for families of sets that appeared in Knuth’s book [13, §7.1.4 Ex. 203,204,236,243].
Table 1: List of operations on family algebra.
Operation Definition Is polytime in DD sizes?
Union $\mathcal{F}\cup\mathcal{G}$$\{S\mid S\in\mathcal{F}\vee S\in\mathcal{G}\}$ Yes [14]
Intersection $\mathcal{F}\cap\mathcal{G}$$\{S\mid S\in\mathcal{F}\wedge S\in\mathcal{G}\}$ Yes [14]
Difference $\mathcal{F}\setminus\mathcal{G}$$\{S\mid S\in\mathcal{F}\wedge S\notin\mathcal{G}\}$ Yes [14]
Symmetric difference $\mathcal{F}\oplus\mathcal{G}$ $(\mathcal{F}\setminus\mathcal{G})\cup(\mathcal{G}\setminus\mathcal{F})$ Yes [14]
Join $\mathcal{F}\sqcup\mathcal{G}$$\{F\cup G\mid F\in\mathcal{F},G\in\mathcal{G}\}$ No (Theorem 3.5)
Disjoint join ˙$\mathcal{F}\operatorname{\dot{\bowtie}}\mathcal{G}$$\{F\cup G\mid F\in\mathcal{F},G\in\mathcal{G},F\cap G=\emptyset\}$ No (Theorem 3.5)
Joint join ^$\mathcal{F}\operatorname{\hat{\bowtie}}\mathcal{G}$$\{F\cup G\mid F\in\mathcal{F},G\in\mathcal{G},F\cap G\neq\emptyset\}$ No (Theorem 3.5)
Meet $\mathcal{F}\sqcap\mathcal{G}$$\{F\cap G\mid F\in\mathcal{F},G\in\mathcal{G}\}$ No (Theorem 3.5)
Delta $\mathcal{F}\operatorname{\boxplus}\mathcal{G}$$\{F\oplus G\mid F\in\mathcal{F},G\in\mathcal{G}\}$ No (Theorem 3.5)
Quotient $\mathcal{F}\operatorname{/}\mathcal{G}$$\{S\mid\forall G\in\mathcal{G}:S\cup G\in\mathcal{F}\wedge S\cap G=\emptyset\}$ No (Theorem 3.9)
Remainder %$\mathcal{F}\operatorname{\%}\mathcal{G}$ $\mathcal{F}\setminus(\mathcal{G}\sqcup(\mathcal{F}\operatorname{/}\mathcal{G}))$ No (Theorem 3.9)
Restrict $\mathcal{F}\bigtriangleup\mathcal{G}$$\{F\in\mathcal{F}\mid\exists G\in\mathcal{G}:G\subseteq F\}$ No (Theorem 3.11)
Permit $\mathcal{F}\oslash\mathcal{G}$$\{F\in\mathcal{F}\mid\exists G\in\mathcal{G}:F\subseteq G\}$ No (Theorem 3.11)
Nonsuperset $\mathcal{F}\operatorname{\searrow}\mathcal{G}$$\{F\in\mathcal{F}\mid\forall G\in\mathcal{G}:G\nsubseteq F\}$ No (Theorem 3.11)
Nonsubset $\mathcal{F}\operatorname{\nearrow}\mathcal{G}$$\{F\in\mathcal{F}\mid\forall G\in\mathcal{G}:F\nsubseteq G\}$ No (Theorem 3.11)
@@
Maximal $\mathcal{F}^{\uparrow}$$\{F\in\mathcal{F}\mid\forall F^{\prime}\in\mathcal{F}:F\subseteq F^{\prime}\Rightarrow F=F^{\prime}\}$ No (Theorem 3.13)
@@
Minimal $\mathcal{F}^{\downarrow}$$\{F\in\mathcal{F}\mid\forall F^{\prime}\in\mathcal{F}:F^{\prime}\subseteq F\Rightarrow F=F^{\prime}\}$ No (Theorem 3.13)
@@
Minimal hitting set $\mathcal{F}^{\sharp}$$\{S\mid\forall F\in\mathcal{F}:S\cap F\neq\emptyset\}^{\downarrow}$ No (Theorem 3.15)
Closure $\mathcal{F}^{\cap}$$\{\bigcap_{S\in\mathcal{F}^{\prime}}S\mid\mathcal{F}^{\prime}\subseteq\mathcal{F}\}$ No (Theorem 3.15)
Previous studies [17, 13] stated that they can be performed in worst-case polynomial time.
@@
Compared to the operations themselves, the time complexity of performing them on ZDDs has not been well investigated. Minato [14] proved that the first four operations in Table 1 can be performed in polynomial time with respect to the size of input ZDDs. However, the complexity of a join operation, the most basic one among the rest, has not been fully clarified. Knuth [13, §7.1.4 Ex. 206] claimed that join, as well as meet and delta, can be performed in worst-case polynomial time, but this claim lacks proof. Conversely, Kawahara et al. [12] suggested that join, as well as disjoint join and joint join, take worst-case exponential time, again without proof. In addition to those reports, Okuno et al. [17] claimed that restrict can be performed in polynomial time, but they used the unproven proposition that join can be performed in polynomial time. Furthermore, Knuth [13, §7.1.4 Ex. 206] stated that the worst-case complexity of the quotient operation was an open problem.
@@ 1.2 Our Contribution
In this paper, we prove that, for the operations in Table 1 aside from the first four operations, there exist polynomial-sized ZDDs such that after taking the operation, the ZDD size becomes exponential. For example, for the join operation, we prove that there exist sequences of families of sets $\{\mathcal{F}_{m}\}$ and $\{\mathcal{G}_{m}\}$ such that the ZDD sizes representing $\mathcal{F}_{m}$ and $\mathcal{G}_{m}$ are polynomial in $m$, while the ZDD size representing $\mathcal{F}_{m}\sqcup\mathcal{G}_{m}$ is exponential in $m$. This result implies that these operations cannot be performed in worst-case polynomial time with respect to the size of input ZDDs. Thus, we refute the statement raised by Knuth [13] and Okuno et al. [17] that join, meet, delta, and restrict can be performed in worst-case polynomial time. We also resolve the worst-case complexity of the quotient operation. Moreover, we also prove that the operations in Table 1, except for the first four operations, cannot be performed in polynomial time even when families are represented by BDDs. Since Table 1 contains all the family algebra operations raised by Knuth [13], this paper concludes what kind of family algebra operations can be performed in polynomial time on BDDs and ZDDs.
@@
Our result is stronger in that the resultant BDD/ZDD’s size remains exponential for any order of elements. BDD/ZDD structures follow a total order of the elements in the base set, and it is known that this element order has a significant impact on the BDD/ZDD size. For example, it is known that a multiplexer function can be represented in linear-sized BDD by managing the ordering while its size becomes exponential when the ordering is terrible [13, p.235]. However, we also prove that for the sequences used in proving the above, the resultant BDD/ZDD’s size is exponential in $m$ regardless of the order of elements. This suggests that we cannot shrink the BDD/ZDD size after taking an operation by managing the element order. Some famous BDD manipulation packages such as CUDD [18] implemented dynamic reordering, the reordering of elements after executing operations to shrink the BDD/ZDD size and thus increase the efficiency of BDD/ZDD manipulations. Nevertheless, our results suggest that the worst-case complexity of carrying out operations cannot be polynomial, even if we employ dynamic reordering.
@@
Note that this follows the research line of Bollig [1] as follows. Yoshinaka et al. [20] refuted Bryant’s conjecture, which is about the complexity of performing operations on BDDs, but their counterexample was somewhat weak in that the order of elements they used was unfavorable for BDD representations. Bollig [1] later resolved this issue by proposing simpler counterexamples. Similar to this, our results imply that the exponential blow-up in taking an operation on BDDs/ZDDs occurs not only when the order of elements is unfavorable but also when it is good for BDD/ZDD representations.
From the viewpoint of applications, BDDs/ZDDs are usually built by applying multiple family algebra operations in combination with some direct construction methods such as Simpath [13] and frontier-based search [11], which are fixed-parameter tractable algorithms with pathwidth. However, the number of required operations stays constant in many applications. If every operation can be performed in polynomial time, we can enjoy the polynomial time complexity in BDD/ZDD sizes even for these applications. However, our results suggest this is not the case except for the first four operations. In addition, although we rely on specific input examples to prove non-polynomial lower bounds, we later discuss that such blow-up may occur for other input; the detailed discussions are in Section 3.5. Therefore, our theoretical results have practical importance.
@@ 2.1 Zero-suppressed Binary Decision Diagram
A zero-suppressed binary decision diagram (ZDD[14] is a rooted directed acyclic graph (DAG)-shaped data structure for representing a family of sets. First, we describe the structure of ZDD. ZDD $\mathtt{Z}$ consists of node set $\mathtt{N}$ and arc set $\mathtt{A}$, where the node set contains terminal nodes $\top$,$\bot$ , ⊥ and other internal nodes. Terminal nodes have no outgoing arcs, while every internal node has two outgoing arcs called lo-arc and hi-arc. The nodes pointed by the lo-arc and the hi-arc outgoing from a node $\mathtt{n}$ are called lo-child $\mathsf{lo}(\mathtt{n})$ and hi-child $\mathsf{hi}(\mathtt{n})$ of $\mathtt{n}$. Every internal node $\mathtt{n}$ is associated with an element called label that is denoted by $\mathsf{lb}(\mathtt{n})$. ZDDs must follow the ordered property: Given a total order of elements $<$, the label of the parent node must precede that of the child node, i.e., $\mathsf{lb}(\mathtt{n})<\mathsf{lb}(\mathsf{lo}(\mathtt{n}))$ and $\mathsf{lb}(\mathtt{n})<\mathsf{lb}(\mathsf{hi}(\mathtt{n}))$ must hold for every internal node $\mathtt{n}$. Note that the child node is always allowed to be a terminal node. Finally, the size of a ZDD is defined by its number of nodes.
Next, we describe the semantics of ZDD.
Definition 2.1.
@@
We briefly compare ZDDs with BDDs. BDD [2] has the same structure (syntax) as ZDD, although its semantics is slightly different. BDDs also follow the ordered property and have the smallest canonical form called reduced BDD. Given the total order $<$ of elements, the size of the reduced BDD of the family $\mathcal{F}$ is denoted by $B_{<}(\mathcal{F})$. The following is a famous result.
Lemma 2.2 ([13, Eq. (126)]).
For any family $\mathcal{F}$ of subsets of a set of $n$ elements and any order $<$ of elements, $B_{<}(\mathcal{F})=O(nZ_{<}(\mathcal{F}))$ and $Z_{<}(\mathcal{F})=O(nB_{<}(\mathcal{F}))$.